# DPImageBench: A Unified Benchmark for Differentially Private Image Synthesis

### v2

@@ Appendix B DP Image Synthesizer Details
B.1 Implementations
DP-MERF: https://github.com/ParkLabML/DP-MERF
DP-Kernel: https://github.com/dihjiang/DP-kernel
DP-LoRA: https://github.com/EzzzLi/DP-LORA
PrivImage: https://github.com/SunnierLee/DP-ImaGen
B.2 Hyper-parameter Settings
B.2.1 Investigated Synthesis Algorithms
This section elaborates on the hyper-parameter settings for DP image synthesis methods. Table XV presents the learning rate and batch size, which are all the same on different datasets for each method. Table XVI presents the number of iterations for finetuning on sensitive datasets. We can calculate the number of iterations as .
@@ B.2.1 Investigated Synthesis Algorithms
PE. Refer to the practical adoption in the original paper [5], we use the pretrained model “imagenet64_cond_270M_250K.pt” as the foundational API333https://github.com/openai/improved-diffusion. For generating images, we use the pretrained model “128x128_diffusion.pt” as the foundational API444https://github.com/openai/guided-diffusion. These APIs are only trained on ImageNet.
DP-GAN. We use BigGAN [82] as the synthesizer in our implementation of DP-GAN, resulting in improved synthesis performance compared to the original DP-GAN [30].
DP-LoRA. The pretraining stage is the same as DP-LDM. During finetuning, we use the LoRA [55] technique to finetune the label embedding layer and attention modules of DM.
@@
B.2.2 Network Architecture
For a fair comparison, we implement the GAN-based methods using the same network architecture of BigGAN generator [82] except for GS-WGAN for its poor performance. Specifically, the generator has a noise vector dimension of 60, a shared feature dimension of 128, and a convolutional feature dimension of 60. For DPDM, PDP-Diffusion, DP-LDM (SD), and PrivImage, we use the same UNet as the generator with a feature dimension of 32, a feature multiplicity of , an attention resolution of . For DP-LDM and DP-LoRA, we use the same UNet as the generator with a feature dimension of 32, a feature multiplicity of , and an attention resolution of . The autoencoder of VAE has a feature dimension of 128, a feature multiplicity of , and an attention resolution of .
@@ B.2.3 Pretraining
We base our implementations on the practical code available in the repositories provided by the original papers. In the default setting of DPImageBench, we use conditional pretraining. Table XVII presents the training iterations used for pretraining DP synthesizers across different methods. For methods that do not use a public dataset, we use their default training parameters for pretraining. For methods using a public dataset, we pretrain these methods for 160 epochs with batch size 1,024. Therefore, the pretrain iterations are (number of ImageNet images) . An epoch refers to one complete pass through the entire dataset, while an iteration represents a single training step on a mini-batch. Consequently, for different datasets, it is expected to have a varying number of iterations per epoch due to differences in dataset size. Table XVII presents the hyper-parameters used for pretraining.
B.2.4 Classifier for Utility Evaluation
@@
B.2.5 Metrics for Fidelity Evaluation
B.2.6 Privacy Cost for Hyper-Parameter Tuning
Current DP image synthesis methods overlook the privacy cost associated with hyper-parameter tuning [5, 32, 4, 31]. DPImageBench follows them and similarly does not account for this cost. To mitigate the impact of neglecting privacy loss during hyper-parameter tuning, we should minimize the hyper-parameter tuning process as much as possible. Therefore, we follow the default hyper-parameters the same as that released in the repository as introduced in Section B. Besides, different synthesis methods often use the same value for shared hyper-parameters when applied to the same synthesis task. The results could be further improved with optimized hyper-parameters beyond those we used.
@@ Appendix C Details Descriptions of Datasets
We evaluate DPImageBench on seven sensitive datasets: (1) MNIST [64], (2) FashionMNIST [65], (3) CIFAR-10 [66], (4) CIFAR-100 [66], (5) EuroSAT [67], (6) CelebA [43], and (7) Camelyon [42]. Besides, the ImageNet ISLVRC2012 [59], MS COCO, and LAION [77] datasets are considered a public dataset. This section introduces the investigated dataset as follows.
ImageNet ISLVRC2012: ImageNet Large Scale Visual Recognition Challenge (ISLVRC) 2012 contains over 14 million images labeled across 1,000 object categories, including animals, vehicles, and others. Each image has a resolution of 256x256 pixels. We use a subset of the whole ImageNet dataset for pretraining like previous works [32, 4]. This subset of the dataset is specifically used for the ImageNet Large Scale Visual Recognition Challenge.
Appendix D Additional Experiments Analysis
@@
Fidelity Evaluations of Camelyon Dataset. Table XVIII shows the fidelity results for various methods on the Camelyon dataset. PE achieves the highest IS score of 4.58 among the studied methods. However, the downstream task accuracy achieved by PE is only 64.9%. These results indicate that although PE can generate images with high practical quality, they are not similar to sensitive images. The precision and recall values of 0.01 and 0.77 further support this perspective. Besides, the FLD of synthetic images achieved by PDP-Diffusion and PrivImage are both negative, i.e., -5.9 and -5.09, suggesting that, according to this metric, the quality of the synthetic images surpasses that of the original images. Additionally, the accuracy achieved by PDP-Diffusion and PrivImage is 84.8% and 87.0%, respectively, which closely approaches the accuracy obtained on the original dataset, i.e., 87.7%.
The Impact of Pretraining. Figure 9 presents the Acc and FLD evaluations of synthetic images under for the FashionMNIST dataset.
@@
Computational Resources. All methods are implemented on a server equipped with four NVIDIA GeForce A6000 Ada GPUs and 512GB of memory. The synthesizer sizes for the GAN-based and diffusion-based methods are 5.6M and 3.8M. For GAN, the generator size is 3.8M, and the discriminator size is 1.8M. Consistently shared hyper-parameters are used across all methods, as outlined in Appendix B.2.1.
@@
Table XIX provides an analysis of GPU memory usage and runtime for the methods studied in DPImageBench using the CIFAR-10 dataset. Overall, the diffusion-based synthesizer requires more GPU resources and has a longer runtime compared to the GAN-based methods. In fact, the computation resource cost heavily depends on multiple factors, like the size of the synthesizer parameters, the batch size, etc. We recommend using the largest possible batch size for finetuning the synthesizers, while it needs more computation resources. However, with the same privacy budget, using a larger batch size results in a smaller scale of DP noise being added during finetuning for each individual image [52], which can lead to improved synthesis results.
@@ Appendix E Error in DP-Promise
DP-Promise [24] proposes reducing the noise introduced by DP during training for highly efficient training. We argue that the DP-Promise does not strictly satisfy DP. In DMs, as introduced in Section 2.2, the image is gradually corrupted with noise until it becomes random noise, while a neural network is trained to reverse this process and recover the original image. DP-Promise proposes achieving DP by utilizing the noisy process in DM. Specifically, in the forward process of DMs, we add noise to the image through , where 777 is required for both training DMs and generating images. is a fixed hyper-parameter, indicates the added noise, and is the noisy image. They theoretically prove that this process naturally satisfies DP with as the DP noise, and its privacy cost decreases as increases. Thus, when is sufficiently large, we do not need to introduce additional noise to the gradient. However, the analysis of DP-Promise overlooks the fact that, in addition to , the training also has access to the noise (DP noise should be unknown) and the time step to calculate the objective function 4. If and are not added with noise, we can easily reconstruct the original image via (by inverting Eq (3)). Therefore, we believe that DP-Promise violates DP, and we do not implement DP-Promise in DPImageBench.

### v3

@@ Appendix B DP Image Synthesizer Details
B.1. Implementations
DP-MERF: https://github.com/ParkLabML/DP-MERF
DP-Kernel: https://github.com/dihjiang/DP-kernel
DP-LoRA: https://github.com/EzzzLi/DP-LORA
PrivImage: https://github.com/SunnierLee/DP-ImaGen
B.2. Hyper-parameter Settings
B.2.1. Investigated Synthesis Algorithms
This section elaborates on the hyper-parameter settings for DP image synthesis methods. Table 13 presents the learning rate and batch size, which are all the same on different datasets for each method. Table 14 presents the number of iterations for finetuning on sensitive datasets. The number of iterations is: .
PE. Refer to the practical adoption in the original paper (Lin et al., 2024), we use the pretrained model “imagenet64_cond_270M_250K.pt” as the foundational API333https://github.com/openai/improved-diffusion. For generating images, we use the pretrained model “128x128_diffusion.pt” as the foundational API.444https://github.com/openai/guided-diffusion These models are only trained on ImageNet. In principle, PE can use any model. For comparison fairness, in the model size experiments in Section 7.2 and all experiments in Sections 7.3 and 7.4, we use the APIs from the PDP-Diffusion pretrained by us in the same experiments.
@@ B.2.1. Investigated Synthesis Algorithms
DP-GAN. We use BigGAN (Brock et al., 2019) as the synthesizer in our implementation of DP-GAN, resulting in improved synthesis performance compared to the original DP-GAN (Xie et al., 2018).
@@
DP-LoRA. The pretraining stage is the same as DP-LDM. During finetuning, we use the LoRA (Hu et al., 2022) technique to finetune the label embedding layer and attention modules of DM.
@@
B.2.2. Network Architecture
For a fair comparison, we implement the GAN-based methods using the same network architecture of BigGAN generator (Brock et al., 2019) except for GS-WGAN for its poor performance. Specifically, the generator has a noise vector dimension of 60, a shared feature dimension of 128, and a convolutional feature dimension of 60. For DPDM, PDP-Diffusion, DP-LDM (SD), and PrivImage, we use the same UNet as the generator with a feature dimension of 32, a feature multiplicity of , an attention resolution of . For DP-LDM and DP-LoRA, we use the same UNet as the generator with a feature dimension of 32, a feature multiplicity of , and an attention resolution of . The autoencoder of VAE has a feature dimension of 128, a feature multiplicity of , and an attention resolution of .
@@ B.2.3. Pretraining
| MNIST (LeCun et al., 1998) | 70,000 | 28281 | 10 | 55,000 | 5,000 | 10,000 | Classification | Sensitive |
| FashionMNIST (Xiao et al., 2017) | 70,000 | 28281 | 10 | 55,000 | 5,000 | 10,000 | Classification | |
| CIFAR-10 (Krizhevsky, 2009) | 60,000 | 32323 | 10 | 45,000 | 5,000 | 10,000 | Classification | |
| CIFAR-100 (Krizhevsky, 2009) | 60,000 | 32323 | 100 | 45,000 | 5,000 | 10,000 | Classification | |
| EuroSAT (Helber et al., 2019) | 27,000 | 64643 | 10 | 21,000 | 2,000 | 4,000 | Classification | |
| CelebA (Liu et al., 2015) | 182,732 | 1281283 | 2 | 145,064 | 17,706 | 19,962 | Classification | |
| Camelyon (Bándi et al., 2019) | 337,340 | 64643 | 2 | 269,538 | 32,898 | 34,904 | Classification | |
| ImageNet ISLVRC2012 (Deng et al., 2009) | 1,431,167 | 2562563 | 1,000 | 1,281,167 | 50,000 | 100,000 | Pretraining | Public |
| Places365 (Zhou et al., 2017) | 2,141,660 | 2562563 | 365 | 1,803,460 | 17,800 | 320,400 | Pretraining |
We base our implementations on the practical code available in the repositories provided by the original papers. In the default setting of DPImageBench, we use conditional pretraining. Table 15 presents the training iterations used for pretraining DP synthesizers across different methods. For methods that do not use a public dataset, we use their default training parameters for pretraining. For methods using a public dataset, we pretrain these methods for 160 epochs with batch size 1,024. Therefore, the pretrain iterations are (number of ImageNet images) . An epoch refers to one complete pass through the entire dataset, while an iteration represents a single training step on a mini-batch. Consequently, for different datasets, it is expected to have a varying number of iterations per epoch due to differences in dataset size. Table 15 presents the hyper-parameters used for pretraining.
B.2.4. Classifier for Utility Evaluation
@@
B.2.5. Metrics for Fidelity Evaluation
We briefly outline the selected fidelity metrics as follows.
FID: FID (Heusel et al., 2017) first uses Inception-v3 (Szegedy et al., 2016) to extract features from synthetic and sensitive images. It estimates multivariate Gaussian distributions from these features and calculates the FID score by measuring the Fréchet distance (Heusel et al., 2017) between the two distributions. A lower FID indicates that the generated images are more similar to the real dataset.
IS: The IS uses the Inception v3 (Szegedy et al., 2016) network to classify each generated image into predefined classes. A higher IS indicates that the synthetic images are both realistic (the model classifies them with high confidence) and diverse (the images span various classes).
Precision and Recall: We begin by extracting features from both sensitive and synthetic images using the Inception v3. Next, for each synthetic image, we identify the closest sensitive image in the feature space. If the distance between the synthetic image and the sensitive image falls below a specified threshold, we classify it as a “true positive.” Precision is then calculated as . For recall, we locate the nearest synthetic image for each sensitive image. Here, a “True Positive” is defined as a pair where the distance between the images is below the threshold, and Recall is calculated as . Achieving higher Precision and Recall is preferable, although there is often a trade-off between them (Kynkäänniemi et al., 2019).
@@ B.2.5. Metrics for Fidelity Evaluation
FLD: We refer to the repository released by Jiralerspong et al. (Jiralerspong et al., 2023). A lower FLD value is better, with negative values indicating the synthetic dataset outperforms the original.
B.2.6. Privacy Cost for Hyper-Parameter Tuning
Current DP image synthesis methods overlook the privacy cost associated with hyper-parameter tuning (Lin et al., 2024; Ghalebikesabi et al., 2023; Li et al., 2024; Dockhorn et al., 2023). DPImageBench follows them and similarly does not account for this cost. To mitigate the impact of neglecting privacy loss during hyper-parameter tuning, we should minimize the hyper-parameter tuning process as much as possible. Therefore, we follow the default hyper-parameters the same as that released in the repository as introduced in Section B. Besides, different synthesis methods often use the same value for shared hyper-parameters when applied to the same synthesis task. The results could be further improved with optimized hyper-parameters beyond those we used.
@@ Appendix C Details Descriptions of Datasets
We evaluate DPImageBench on seven widely studied sensitive datasets, including two gray and five colorful datasets: (1) MNIST (LeCun et al., 1998), (2) FashionMNIST (Xiao et al., 2017), (3) CIFAR-10 (Krizhevsky, 2009), (4) CIFAR-100 (Krizhevsky, 2009), (5) EuroSAT (Helber et al., 2019), (6) CelebA (Liu et al., 2015), and (7) Camelyon (Bándi et al., 2019). Besides, the ImageNet ISLVRC2012 (Deng et al., 2009) and Places365 (Zhou et al., 2017) datasets are considered a public dataset. This section introduces the investigated dataset as follows.
@@
ImageNet ISLVRC2012: ImageNet Large Scale Visual Recognition Challenge (ISLVRC) 2012 contains over 14 million images labeled across 1,000 object categories, including animals, vehicles, and others. Each image has a resolution of 256x256 pixels. We use a subset of the whole ImageNet dataset for pretraining like previous works (Ghalebikesabi et al., 2023; Li et al., 2024). This subset of the dataset is specifically used for the ImageNet Large Scale Visual Recognition Challenge.
Appendix D Additional Experiments Analysis
D.1. Fidelity Evaluations of Camelyon Dataset
@@
Table 17 shows the fidelity results for various methods on the Camelyon dataset. PE achieves the highest IS score of 4.58 among the studied methods. However, the downstream task accuracy achieved by PE is only 64.9%. These results indicate that although PE can generate images with high practical quality, they are not similar to sensitive images. The precision and recall values of 0.01 and 0.77 further support this perspective. Besides, the FLD of synthetic images achieved by PDP-Diffusion and PrivImage are both negative, i.e., -5.9 and -5.09, suggesting that, according to this metric, the quality of the synthetic images surpasses that of the original images. Additionally, the accuracy achieved by PDP-Diffusion and PrivImage is 84.8% and 87.0%, respectively, which closely approaches the accuracy obtained on the original dataset, i.e., 87.7%.
D.2. Computational Resources
@@
All methods are implemented on a server equipped with four NVIDIA GeForce A6000 Ada GPUs and 512GB of memory. The synthesizer sizes for the GAN-based and diffusion-based methods are 5.6M and 3.8M. For GAN, the generator size is 3.8M, and the discriminator size is 1.8M. Consistently shared hyper-parameters are used across all methods, as outlined in Appendix B.2.1.
| Batch size | M | M | M | M |
| 4,096 | 78.4 | 78.2 | 76.0 | 79.5 |
| 16,384 | 78.5 | 78.6 | 78.8 | 79.5 |
@@
Table 18 provides an analysis of GPU memory usage and runtime for the methods studied in DPImageBench using the CIFAR-10 dataset. Overall, the diffusion-based synthesizer requires more GPU resources and has a longer runtime compared to the GAN-based methods. In fact, the computation resource cost heavily depends on multiple factors, like the size of the synthesizer parameters, the batch size, etc. We recommend using the largest possible batch size for finetuning the synthesizers, while it needs more computation resources. However, with the same privacy budget, using a larger batch size results in a smaller scale of DP noise being added during finetuning for each individual image (Dwork et al., 2014), which can lead to improved synthesis results.
D.3. Visualize Results
Figure 9 illustrates the examples of synthetic images under .
D.4. Impact on Batch Size of DP-SGD
Table 19 presents the accuracy (Acc) of synthetic images generated by PrivImage under a privacy budget of , using the sensitive dataset CIFAR-10. The table examines the effect of batch sizes—4,096 and 16,384 (the latter being the default batch size in PrivImage (Li et al., 2024))—across four synthesizer model sizes of 3.8M, 11.1M, 19.6M, and 44.2M parameters. For the batch size of 4,096, accuracy varies from 76.0% (at 19.6M) to 79.5% (at 44.2M), indicating a slight upward trend as model size increases, though a dip occurs at 19.6M. In contrast, with the larger batch size of 16,384, accuracy consistently improves across model sizes, ranging from 78.5% (at 3.8M) to 79.5% (at 44.2M), with a peak of 78.8% at 19.6M. These findings suggest that larger batch sizes in DP-SGD significantly enhance performance, particularly for larger model sizes, likely because they reduce the noise scale in gradient updates, leading to more stable training and better convergence under differential privacy constraints.
D.5. The Impact of Pretraining
Figure 8 presents the Acc and FLD evaluations of synthetic images under for the FashionMNIST dataset.
Table 20 presents the accuracy (Acc) and Fréchet Latent Distance (FLD) of synthetic images generated by PDP-Diffusion and PrivImage under a privacy budget of . The sensitive dataset MNIST and public dataset ImageNet are used, with pretraining iterations ranging from 25K to 400K. Results show that both methods achieve high Acc (above 97%), with PrivImage generally exhibiting lower FLD values, indicating better distributional similarity to the sensitive data across varying pretraining steps.
Figure 11 compares the Acc and FLD for the sensitive datasets FashionMNIST when pretraining with public datasets ImageNet and Places365. For FashionMNIST, the average accuracy scores with ImageNet and Places365 are 77.2 and 76.1. Pretraining on ImageNet consistently outperforms Places365, except in the case of PDP-Diffusion trained on FashionMNIST, where the difference is marginal (0.3%).
| PDP-Diffusion | 25K | 50K | 100K | 200K | 400K |
| Acc (%) | 97.4 | 97.6 | 97.4 | 97.4 | 97.1 |
| FLD | 2.48 | 2.56 | 2.83 | 3.40 | 3.00 |
| PrivImage | 25K | 50K | 100K | 200K | 400K |
| Acc (%) | 97.8 | 97.4 | 97.5 | 97.8 | 97.2 |
| FLD | 2.23 | 2.48 | 2.62 | 2.80 | 2.90 |
D.6. Combining DP-FETA with other methods
We incorporate a shortcut process into the DP training, extending it to other methods. Specifically, we extract central images from sensitive datasets and train synthesizers on these images prior to training on the sensitive images. Central images are perturbed with Gaussian noise for DP. PE is a fine-tuning-free method, so we exclude it from this analysis. Table 21 shows the results of integrating the DP training shortcut into other methods.
| Algorithm | With DP Training Shortcuts | Original | ||
| (CIFAR-10) | Acc (%) | FLD | Acc (%) | FLD |
| DP-MERF | 57.7 | 24.5 | 62.2 | 29.2 |
| DP-NTK | 70.6 | 31.0 | 76.3 | 36.4 |
| DP-Kernel | 78.0 | 18.5 | 70.0 | 21.3 |
| GS-WGAN | 58.9 | 27.2 | 56.7 | 28.1 |
| DP-GAN | 55.4 | 22.8 | 70.3 | 23.9 |
| DPDM | 87.3 | 4.6 | 85.6 | 6.6 |
| PDP-Diffusion | 85.0 | 4.7 | 85.4 | 4.9 |
| DP-LDM (SD) | 80.6 | 13.3 | 81.6 | 11.7 |
| DP-LDM | 85.3 | 18.8 | 86.3 | 15.4 |
| DP-LoRA | 86.3 | 14.2 | 83.8 | 14.8 |
| PrivImage | 86.4 | 4.6 | 87.1 | 4.3 |
| Average | 75.6 | 16.7 | 76.8 | 17.9 |
Building DP training shortcut may damage the synthetic performance. Table 21 shows that incorporating a DP training shortcut may impair the synthetic performance of methods with pretraining. For instance, the Acc and FLD of DP-LDM (SD) decline from 81.6% and 11.7 to 80.6% and 13.3 when shortcuts are added. This is because the shortcut includes training on the central images after pretraining, which may damage the pretraining synthesizer. For methods without pretraining, the fidelity of synthetic images (measured by FLD) is consistently improved, while the improvement of utility (measured by Acc) fluctuates. These results show that “,” indicating that a straightforward algorithmic combination does not always improve synthetic performance, and highlighting the need for further research into effectively integrating the strengths of diverse algorithms.
| FLD | 5% | 5% (rand.) | 5%+5% (rand.) | 10% (rand.) | Whole |
| DP-MERF | 30.2 | 31.6 | 29.8 | 29.7 | 28.4 |
| DP-NTK | 50.0 | 50.5 | 55.9 | 53.6 | 49.8 |
| DP-Kernel | 31.5 | 30.1 | 30.3 | 31.1 | 30.0 |
| DP-GAN | 18.9 | 18.9 | 21.8 | 21.6 | 29.5 |
| PE | 4.5 | 6.5 | 12.2 | 10.6 | 8.5 |
| DP-FETA | 10.7 | 12.4 | 9.3 | 12.8 | 10.6 |
| PDP-Diffusion | 4.7 | 10.7 | 6.1 | 9.3 | 7.2 |
| DP-LDM (SD) | 5.1 | 11.7 | 7.2 | 10.6 | 9.0 |
| DP-LDM | 12.2 | 15.2 | 13.0 | 14.4 | 14.1 |
| DP-LoRA | 8.1 | 11.5 | 8.0 | 11.4 | 9.3 |
@@ Appendix E Error in DP-Promise
DP-Promise (Wang et al., 2024) proposes reducing the noise introduced by DP during training for highly efficient training. We argue that the DP-Promise does not strictly satisfy DP. In DMs, as introduced in Section 2.2, the image is gradually corrupted with noise until it becomes random noise, while a neural network is trained to reverse this process and recover the original image. DP-Promise proposes achieving DP by utilizing the noisy process in DM. Specifically, in the forward process of DMs, we add noise to the image through , where 777 is required for both training DMs and generating images. is a fixed hyper-parameter, indicates the added noise, and is the noisy image. They theoretically prove that this process naturally satisfies DP with as the DP noise, and its privacy cost decreases as increases. Thus, when is sufficiently large, we do not need to introduce additional noise to the gradient. However, the analysis of DP-Promise overlooks the fact that, in addition to , the training also has access to the noise (DP noise should be unknown) and the time step to calculate the objective function 4. If and are not added with noise, we can easily reconstruct the original image via (by inverting Eq (3)). Therefore, we believe that DP-Promise violates DP, and we do not implement DP-Promise in DPImageBench.